OpenStudy (anonymous):

How to measure the Co2 in alka-seltzer and then calculate the 3NaHCO3? How do you measure the mass in grams of NaHCO3, that is found in each tablet of Alka-Seltzer?

OpenStudy (abb0t):

You can start by writing down the equation of your reaction with. Then, write the mass of the alka-seltzer, which can be found in the packet or by using a weight measuring machine.

OpenStudy (abb0t):

Don't forget to balance the equation, that way, you can use dimensional analysis to calculate the theoretical yield of what you want to find.

OpenStudy (anonymous):

C6H8O7+3NaHCO3----> 3H2O+3CO2+Na3C6H5O7

OpenStudy (anonymous):

Weight is 3.241 grams

OpenStudy (abb0t):

Now, find the moles of sodium carbonate.s

OpenStudy (abb0t):

Do you know how to do that?

OpenStudy (anonymous):

no i dont

OpenStudy (abb0t):

Well, start with the given grams of alkazelteser that you have. To start, compute the molar mass of it, which you can do by finding the sum of it's molecular formula. For example, molar mass has the units of \(\frac{mass}{mole}\), therefore, if you have H\(_2\), the molar mass of that is 1.008 \(\times\) 2 = 2.016 \(\frac{grams}{mol}\). Does that make sense? And very similar to how you did in algebra. you are going to use that relationship to cancel out the units. given grams of Alka-Seltz \(\times\) \(\frac{mol}{mass~of~alka-setlz}\) notice that the units of grams of alka-seltz cancel out? You will be left with moles. Then, take: mol of Alka-seltz \(\times\) \(\frac{\sf \color{red}{3}mol~NaHCO_3}{\sf \color{blue}{1} mol~alka-seltz}\) again, note that the units of mole cancel out. Now you have Mol's of NaHCO\(_3\)! Finally, multiply by the molar mass of NaHCO\(_3\) to get the molar mass! Now, based on your equation, it tells you that you have a 1:1 ratio of NaHCO\(_3\). So, mole of NaHCO\(_3\) = mol of CO\(_2\). Now, this is the amount of CO\(_2\) you would get if All the NaHCO\(_3\) fully reacted. It is called the theoretical yield of CO\(_2\). It is the amount of CO\(_2\) you are expecting to get, assuming that all the NaHCO\(_3\) reacts. Convert it to mass, SIMILARLY to how you did above.

OpenStudy (abb0t):

Do you follow me? Or do you need more help. i am here.

OpenStudy (anonymous):

yes sort of, so woukd the mol be 0.0073

OpenStudy (anonymous):

Molar mass of NaHCO3 is 84g

OpenStudy (anonymous):

I don't fully understand how the last part

OpenStudy (abb0t):

I have to step out for a minute, but if you follow my steps, you should get the right answer. Sorry.

OpenStudy (anonymous):

Alright thank you!